# On the Expressiveness of Visual Prompt Experts

### v5

@@ Abstract
Visual Prompt Tuning (VPT) has proven effective for parameter-efficient adaptation of pre-trained vision models to downstream tasks by inserting task-specific learnable prompt tokens. Despite its empirical success, a comprehensive theoretical understanding of VPT remains an active area of research. Building on the recently established connection between Mixture of Experts (MoE) and prompt-based methods, wherein each attention head can be conceptualized as a composition of multiple MoE models, we reinterpret VPT as the introduction of new prompt experts into these MoE structures. We identify a key limitation in existing VPT frameworks: the restricted functional expressiveness of prompt experts, which remain static and thus limited in their adaptability. To address this, we propose Visual Adaptive Prompt Tuning (VAPT), a novel method that endows prompt experts with enhanced expressiveness while preserving parameter efficiency. Empirical evaluations on VTAB-1K and FGVC demonstrate that VAPT achieves substantial performance improvements, surpassing fully fine-tuned baselines by 7.34% and 1.04%, respectively. Moreover, VAPT consistently outperforms VPT while requiring fewer additional parameters. Furthermore, our theoretical analysis indicates that VAPT achieves optimal sample efficiency. Collectively, these results underscore the theoretical grounding and empirical advantages of our approach.
@@ 1 Introduction
Foundational vision models, pre-trained on large-scale datasets dosovitskiy2020image ; radford2021learning ; kirillov2023segment , have demonstrated remarkable success and robust generalization across a wide range of computer vision tasks. As a result, fine-tuning these models for specific downstream tasks has become a widely adopted paradigm iofinova2022well . However, fully fine-tuning large foundational models can be computationally prohibitive, leading to growing interest in parameter-efficient fine-tuning (PEFT) techniques cai2020tinytl ; zhang2020side ; hu2021lora , which update only a small subset of parameters. Among these methods, Visual Prompt Tuning (VPT) jia2022visual has emerged as a simple yet powerful approach that appends learnable prompt tokens to the input, which serve as task-specific instructions to guide the pre-trained transformer model. Despite VPT’s empirical effectiveness, its theoretical underpinnings remain an active area of research petrov2023prompting ; oymak2023role ; wang2024universality .
@@
Recently, le2024mixture established a formal connection between attention mechanisms vaswani2017attention , prompt-based methods, and Mixture of Experts (MoE) models Jacob_Jordan-1991 ; Quoc-conf-2017 , yielding new insights into the design and optimization of prompting strategies. Their analysis demonstrates that each attention head in a transformer can be equivalently interpreted as a composition of multiple MoE models stacked together. Within this framework, VPT corresponds to fine-tuning these implicit, pre-trained MoE structures by introducing new, learnable prompt experts. These prompt experts collaborate with the pre-trained experts to facilitate effective task adaptation. This connection opens avenues for deeper theoretical investigations and the development of advanced strategies for prompt-based learning le2024revisiting ; le2024adaptive ; diep2025zero .
@@
Building on this MoE interpretation, we identify a key limitation in the standard formulation of VPT. While the pre-trained experts within attention heads are linear functions of the input features, the newly introduced prompt experts are modeled as constant, input-invariant vectors. Given that effective adaptation relies on collaboration between these prompt and pre-trained experts, we hypothesize that this functional disparity, specifically, the restricted expressiveness of static prompts, may constrain VPT’s adaptation efficacy. Furthermore, this design deviates from standard MoE practices, where experts are typically designed to be input-adaptive. This observation naturally raises an important question about the potential of adaptive prompts, wherein prompts are modeled as more expressive functions of the input. Although enhancing adaptivity might raise concerns about parameter overhead, we propose a novel formulation that injects input-dependent prompts while preserving the parameter efficiency characteristic of VPT. We term this approach Visual Adaptive Prompt Tuning (VAPT).
@@
The VAPT design comprises two main components: token-wise projectors and a shared feature projector, which leverage global information from input features to generate adaptive prompt tokens. A key advantage of VAPT is its efficiency: it achieves enhanced expressiveness with minimal computational overhead, adding only 0.6% FLOPs relative to VPT and requiring fewer trainable parameters. Additionally, the token-wise projectors and channel-wise convolutions result in a structurally simple formulation of the prompt experts. This simplicity enables a rigorous theoretical analysis (see Section 5), a contribution largely absent in the prior prompt-tuning literature. Our analysis demonstrates that VAPT attains optimal sample efficiency for prompt estimation. Empirical results strongly corroborate these theoretical findings. For example, in a low-data regime on the Stanford Dogs dataset khosla2011novel (using only 1% of training data), VAPT achieves 60.1% accuracy, a stark contrast to VPT’s 3.6%. Moreover, on standard benchmarks VTAB-1K zhai2019large and FGVC jia2022visual , VAPT significantly improves performance over fully fine-tuned baselines by 7.34% and 1.04%, respectively. Crucially, VAPT consistently outperforms VPT across benchmarks despite utilizing fewer trainable parameters. These results underscore the theoretical soundness and empirical effectiveness of VAPT.
@@
Contributions. 1. From MoE perspective, we identify a key limitation in the formulation of VPT: its prompt experts are input-invariant, thereby constraining their functional expressiveness. 2. We introduce VAPT, a novel formulation that injects input-adaptive prompt experts while preserving the parameter-efficiency of VPT. 3. VAPT’s simple formulation enables a theoretical analysis to validate its effectiveness, an aspect largely missing in prior work. Our theoretical analysis demonstrates that VAPT attains optimal sample efficiency for prompt estimation, providing a rigorous foundation for its practical value. 4. Extensive experiments show that VAPT consistently surpasses VPT with fewer trainable parameters, validating its effectiveness and efficiency both theoretically and empirically.
Organization. Section 2 reviews visual prompt tuning and mixture of experts. In Section 3, we discuss the connection between VPT and MoE, highlighting the limited expressiveness of the standard VPT formulation. Section 4 details our proposed method, designed to achieve higher accuracy with greater parameter efficiency. We present our theoretical analysis in Section 5, followed by empirical results in Section 6. Finally, Section 7 concludes the paper and suggests potential future directions.
@@
Notation. For $n\in\mathbb{N}$, let $[n]=\{1,2,\ldots,n\}$. For a set $S$, $|S|$ denotes its cardinality. Given a vector $u=(u_{1},u_{2},\ldots,u_{d})\in\mathbb{R}^{d}$ and $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}^{d}$, define $u^{\alpha}=u_{1}^{\alpha_{1}}u_{2}^{\alpha_{2}}\ldots u_{d}^{\alpha_{d}}$, $|u|=u_{1}+u_{2}+\ldots+u_{d}$ and $\alpha!=\alpha_{1}!\alpha_{2}!\ldots\alpha_{d}!$. Let $\|u\|$ denote the Euclidean norm of $u$. For positive sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}$, we write $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for all $n\in\mathbb{N}$ for some $C>0$. The notation $a_{n}=\mathcal{O}_{P}(b_{n})$ indicates $a_{n}/b_{n}$ is stochastically bounded.
@@ 2.1 Visual Prompt Tuning
Vision Transformer (ViT) dosovitskiy2020image has proven to be a powerful backbone architecture for visual recognition. A ViT contains $L$ blocks, each comprising a multi-head self-attention (MSA) layer followed by a feed-forward network (FFN). For clarity, we consider the $l$-th ViT block. Let $\tilde{{\bm{X}}}^{(l)}=\left[{\bm{x}}_{1}^{(l)},\dots,{\bm{x}}_{N}^{(l)}\right]^{\top}\in\mathbb{R}^{N\times d}$ be the input tokens, where $N$ is the sequence length, and $d$ is the embedding dimension. The MSA layer processes ${\bm{X}}^{Q}={\bm{X}}^{K}={\bm{X}}^{V}=\tilde{{\bm{X}}}^{(l)}\in\mathbb{R}^{N\times d}$ as queries, keys, and values, producing:
@@
Visual Prompt Tuning (VPT) jia2022visual adapts a pre-trained ViT by appending $N_{p}$ learnable prompt parameters ${\bm{P}}^{(l)}=\left[\bm{p}_{1}^{(l)},\dots,\bm{p}_{N_{p}}^{(l)}\right]^{\top}\in\mathbb{R}^{N_{p}\times d}$, (where $N_{p}$ is the prompt length), to the input of each ViT block, thereby modifying the MSA layer’s output as follows:
@@ 2.2 Mixture of Experts
The Mixture of Experts (MoE) is a class of statistical machine learning frameworks that combines multiple models, known as experts, to produce more expressive and accurate predictions Jacob_Jordan-1991 ; jordan1994hierarchical . An MoE model typically consists of $N^{\prime}$ expert functions, $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{v}}$ for $i\in[N^{\prime}]$, along with a gating function, $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N^{\prime}}$ which assigns weights to the experts based on learned score functions, $s_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. For a given input ${\bm{h}}\in\mathbb{R}^{d}$, the MoE model generates its output as:
@@
where $G({\bm{h}})=\mathrm{softmax}(s_{1}({\bm{h}}),\ldots,s_{N^{\prime}}({\bm{h}}))$. Recent studies reveal connections between the attention mechanism, prompt-based methods, and MoE frameworks le2024mixture ; le2024revisiting , presenting new promising opportunities to investigate prompt-based techniques through MoE lens.
@@ 3 Motivation
Mixture of Experts meets Visual Prompt Tuning. We begin by discussing the connection between VPT and MoE. Adopting the notation from Section 2.1 and referring to Equation (2), let ${\bm{X}}^{(l)}=\left[{{\bm{x}}_{1}^{(l)}}^{\top},\dots,{{\bm{x}}_{N}^{(l)}}^{\top}\right]^{\top}\in\mathbb{R}^{Nd}$ be the concatenation of all $N$ input tokens to the $l$-th MSA layer. For notational simplicity, the layer superscript $(l)$ is omitted in the subsequent derivations. As shown in recent work le2024mixture , each output vector $\tilde{{\bm{h}}}_{m,i}$ within the $m$-th attention head can be equivalently expressed as the output of an MoE model with input ${\bm{X}}$. Specifically, we define the set of experts as:
@@
These formulations reveal that each attention head in ViT implicitly encodes multiple MoE models, $\tilde{{\bm{h}}}_{m,i}$ for $i\in[N]$. This contrasts with conventional MoE layers Quoc-conf-2017 , where experts and their gating functions typically operate on individual token embeddings ${\bm{x}}_{i}$. In our formulation, the expert networks and their score functions process the entire input sequence ${\bm{X}}$. Furthermore, these initial experts $f_{1},\dots,f_{N}$ and score functions are inherently part of the pre-trained ViT, with their parameters embedded within the model weights (see Equations (3) and (5)), thereby constituting pre-trained experts. Meanwhile, the new prompt experts $f_{N+1},\dots,f_{N+N_{p}}$, whose learnable parameters are contained within prompts (see Equations (4) and (6)), are introduced to efficiently adapt the model to downstream tasks. Consequently, VPT can be viewed as an efficient method for fine-tuning these implicit MoE models by adding new prompt experts. These added experts effectively act as downstream task experts, enabling specialization for new tasks without retraining the entire model.
Restricted Functional Expressiveness of Prompt Experts. Equation (4) reveals a key limitation of prompt experts $f_{N+1},\dots,f_{N+N_{p}}$. Although they are formally functions of ${\bm{X}}$, they are represented by fixed prompt vectors ${\bm{p}}_{1},\dots,{\bm{p}}_{N_{p}}$ that remain constant regardless of the input. While their associated score functions $s_{i,N+j^{\prime}}$ in Equation (6) are input-dependent linear functions, the functional form realized by the prompt expert itself is static. This departs from typical MoE usage in the literature, where experts are adaptive functions of the input. Furthermore, it contrasts with the pre-trained experts $f_{1},\dots,f_{N}$, which are linear functions of ${\bm{X}}$ (see Equation (3)), making them comparatively more expressive. We hypothesize that this limited flexibility may constrain the effectiveness of VPT as a fine-tuning strategy. Supporting this, prior work petrov2023prompting demonstrates that prompt tuning can only add a bias term to the output of an attention block, thereby restricting representational capacity. This observation naturally motivates the central question: Can the performance of visual prompt tuning be improved by making prompt experts more expressive? Addressing this, we propose a novel prompt formulation to enhance their expressiveness in Section 4.

### v6

# Revisit Visual Prompt Tuning: The Expressiveness of Prompt Experts

@@ Abstract
Visual Prompt Tuning (VPT) has proven effective for parameter-efficient adaptation of pre-trained vision models to downstream tasks by inserting task-specific learnable prompt tokens. Despite its empirical success, a comprehensive theoretical understanding of VPT remains an active area of research. Building on the recently established connection between Mixture of Experts (MoE) and prompt-based methods, wherein each attention head can be conceptualized as a composition of multiple MoE models, we reinterpret VPT as the introduction of new prompt experts into these MoE structures. We identify a key limitation in existing VPT frameworks: the restricted functional expressiveness of prompt experts, which remain static and thus limited in their adaptability. To address this, we propose Visual Adaptive Prompt Tuning (VAPT), a novel method that endows prompt experts with enhanced expressiveness while preserving parameter efficiency. Empirical evaluations on VTAB-1K and FGVC demonstrate that VAPT achieves substantial performance improvements, surpassing fully fine-tuned baselines by 7.34% and 1.04%, respectively. Moreover, VAPT consistently outperforms VPT while requiring fewer additional parameters. Furthermore, our theoretical analysis indicates that VAPT achieves optimal sample efficiency. Collectively, these results underscore the theoretical grounding and empirical advantages of our approach. Our code is publicly available at https://github.com/Minhchuyentoancbn/VAPT.
footnotetext: Equal contribution   Work done while at Qualcommfootnotetext: Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.
@@ 1 Introduction
Foundational vision models, pre-trained on large-scale datasets (Dosovitskiy, 2020; Radford et al., 2021; Kirillov et al., 2023), have demonstrated remarkable success and robust generalization across a wide range of computer vision tasks. As a result, fine-tuning these models for specific downstream tasks has become a widely adopted paradigm (Iofinova et al., 2022). However, fully fine-tuning large foundational models can be computationally prohibitive, leading to growing interest in parameter-efficient fine-tuning (PEFT) techniques (Cai et al., 2020; Zhang et al., 2020; Hu et al., 2021), which update only a small subset of parameters. Among these methods, Visual Prompt Tuning (VPT) (Jia et al., 2022) has emerged as a simple yet powerful approach that appends learnable prompt tokens to the input, which serve as task-specific instructions to guide the pre-trained transformer model. Despite VPT’s empirical effectiveness, its theoretical underpinnings remain an active area of research (Petrov et al., 2023; Oymak et al., 2023; Wang et al., 2024).
@@
Recently, Le et al. (2024) established a formal connection between attention mechanisms (Vaswani, 2017), prompt-based methods, and Mixture of Experts (MoE) models (Jacobs et al., 1991; Shazeer et al., 2017), yielding new insights into the design and optimization of prompting strategies. Their analysis demonstrates that each attention head in a transformer can be equivalently interpreted as a composition of multiple MoE models stacked together. Within this framework, VPT corresponds to fine-tuning these implicit, pre-trained MoE structures by introducing new, learnable prompt experts. These prompt experts collaborate with the pre-trained experts to facilitate effective task adaptation. This connection opens avenues for deeper theoretical investigations and the development of advanced strategies for prompt-based learning (Le et al., 2025b; a; Diep et al., 2025).
@@
Building on this MoE interpretation, we identify a key limitation in the standard formulation of VPT. While the pre-trained experts within attention heads are linear functions of the input features, the newly introduced prompt experts are modeled as constant, input-invariant vectors. Given that effective adaptation relies on collaboration between these prompt and pre-trained experts, we hypothesize that this functional disparity, specifically, the restricted expressiveness of static prompts, may constrain VPT’s adaptation efficacy. This design also deviates from standard MoE practices, where experts are typically designed to be input‑adaptive. However, increasing the expressiveness of prompt experts raises concerns about parameter and computational overhead. This leads us to the following question:
(Q) Can we improve model performance by increasing the expressiveness of prompt experts while maintaining parameter efficiency?
@@
To answer this question, we propose Visual Adaptive Prompt Tuning (VAPT), a novel approach that incorporates input-dependent prompt experts while maintaining the parameter efficiency characteristic of VPT. The VAPT design comprises two main components: token-wise projectors and a shared feature projector, which leverage global information from input features to generate adaptive prompt tokens. A key advantage of VAPT is its efficiency: it achieves enhanced expressiveness with minimal computational overhead, adding only 0.6% FLOPs relative to VPT and requiring fewer trainable parameters. Additionally, the token-wise projectors and channel-wise convolutions result in a structurally simple formulation of the prompt experts. This simplicity enables a rigorous theoretical analysis (see Section 5), a contribution largely absent in the prior prompt-tuning literature. Our analysis demonstrates that VAPT attains optimal sample efficiency for prompt estimation. Empirical results strongly corroborate these theoretical findings. For example, in a low-data regime on the Stanford Dogs dataset (Khosla et al., 2011) (using only 1% of training data), VAPT achieves 60.1% accuracy, a stark contrast to VPT’s 3.6%. Moreover, on standard benchmarks VTAB-1K (Zhai et al., 2019) and FGVC (Jia et al., 2022), VAPT significantly improves performance over fully fine-tuned baselines by 7.34% and 1.04%, respectively. Crucially, VAPT consistently outperforms VPT across benchmarks despite utilizing fewer parameters, underscoring its theoretical and empirical strengths.
@@
Contributions. 1. From MoE perspective, we identify a key limitation in the formulation of VPT: its prompt experts are input‑invariant, thereby constraining their functional expressiveness. 2. We introduce VAPT, a novel formulation that injects input‑adaptive prompt experts while preserving the parameter‑efficiency of VPT. 3. VAPT’s simple formulation enables a theoretical analysis to validate its effectiveness, an aspect largely missing in prior work. Our theoretical analysis demonstrates that VAPT attains optimal sample efficiency for prompt estimation, providing a rigorous foundation for its practical value. 4. Extensive experiments show that VAPT consistently surpasses VPT with fewer trainable parameters, validating its effectiveness and efficiency both theoretically and empirically.
@@
Notation. For $n\in\mathbb{N}$, let $n]=\{1,2,\ldots,n\}$. For a set $S$, $|S|$ denotes its cardinality. Given a vector $u=(u_{1},u_{2},\ldots,u_{d})\in\mathbb{R}^{d}$ and $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}^{d}$, define $u^{\alpha}=u_{1}^{\alpha_{1}}u_{2}^{\alpha_{2}}\ldots u_{d}^{\alpha_{d}}$, $|u|=u_{1}+u_{2}+\ldots+u_{d}$ and $\alpha!=\alpha_{1}!\alpha_{2}!\ldots\alpha_{d}!$. Let $\|u\|$ denote the Euclidean norm of $u$. For positive sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}$, we write $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for all $n\in\mathbb{N}$ for some $C>0$. The notation $a_{n}=\mathcal{O}_{P}(b_{n})$ indicates $a_{n}/b_{n}$ is stochastically bounded.
@@ 2.1 Visual Prompt Tuning
Vision Transformer (ViT) (Dosovitskiy, 2020) has proven to be a powerful backbone architecture for visual recognition. A ViT contains $L$ blocks, each comprising a multi-head self-attention (MSA) layer followed by a feed-forward network (FFN). For clarity, we consider the $l$-th ViT block. Let $\tilde{{\bm{X}}}^{(l)}=\left[{\bm{x}}_{1}^{(l)},\dots,{\bm{x}}_{N}^{(l)}\right]^{\top}\in\mathbb{R}^{N\times d}$ be the input tokens, where $N$ is the sequence length, and $d$ is the embedding dimension. The MSA layer processes ${\bm{X}}^{Q}={\bm{X}}^{K}={\bm{X}}^{V}=\tilde{{\bm{X}}}^{(l)}\in\mathbb{R}^{N\times d}$ as queries, keys, and values, producing:
@@
Visual Prompt Tuning (VPT) (Jia et al., 2022) adapts a pre‑trained ViT by appending $N_{p}$ learnable prompt parameters ${\bm{P}}^{(l)}=\left[\bm{p}_{1}^{(l)},\dots,\bm{p}_{N_{p}}^{(l)}\right]^{\top}\in\mathbb{R}^{N_{p}\times d}$, (where $N_{p}$ is the prompt length), to the input of each ViT block, thereby modifying the MSA layer’s output as follows:
@@ 2.2 Mixture of Experts
Mixture of Experts (MoE) is a class of statistical machine learning frameworks that combines multiple models, known as experts, to produce more expressive and accurate predictions (Jacobs et al., 1991; Jordan & Jacobs, 1994). An MoE model typically consists of $N^{\prime}$ expert functions, $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{v}}$ for $i\in[N^{\prime}]$, along with a gating function, $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N^{\prime}}$ which assigns weights to experts based on learned score functions, $s_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. For a given input ${\bm{h}}\in\mathbb{R}^{d}$, the MoE model generates output as:
@@
where $G({\bm{h}})=\mathrm{softmax}(s_{1}({\bm{h}}),\ldots,s_{N^{\prime}}({\bm{h}}))$. Recent studies reveal connections between the attention mechanism, prompt-based methods, and MoE frameworks (Le et al., 2024; 2025b), presenting new promising opportunities to investigate prompt-based techniques through MoE lens.
@@ 3 Motivation
Mixture of Experts meets Visual Prompt Tuning. We begin by discussing the connection between VPT and MoE. Following the notation from Section 2.1 and Equation (2), let ${\bm{X}}^{(l)}=\left[{{\bm{x}}_{1}^{(l)}}^{\top},\dots,{{\bm{x}}_{N}^{(l)}}^{\top}\right]^{\top}\in\mathbb{R}^{Nd}$ be the concatenation of all $N$ input tokens to the $l$-th MSA layer. For notational simplicity, the layer superscript $(l)$ is omitted in the subsequent derivations. As shown in Le et al. (2024), each output vector $\tilde{{\bm{h}}}_{m,i}$ within the $m$-th attention head can be equivalently expressed as the output of an MoE model with input $\bm{X}$}. Specifically, we define the set of experts as:
@@
These formulations reveal that each attention head in ViT implicitly encodes multiple MoE models, $\tilde{{\bm{h}}}_{m,i}$ for $i\in[N]$. This contrasts with conventional MoE layers (Shazeer et al., 2017), where experts and their gating functions typically operate on individual token embeddings ${\bm{x}}_{i}$. In our formulation, the expert networks and their score functions process the entire input sequence $\bm{X}$}. Furthermore, these initial experts $f_{1},\dots,f_{N}$ and score functions are inherently part of the pre-trained ViT, with their parameters embedded within the model weights (see Equations (3) and (5)), thereby constituting pre-trained experts. Meanwhile, the new prompt experts $f_{N+1},\dots,f_{N+N_{p}}$, whose learnable parameters are contained within prompts (see Equations (4) and (6)), are introduced to efficiently adapt the model to downstream tasks. Consequently, VPT can be viewed as an efficient method for fine-tuning these implicit MoE models by adding new prompt experts. These added experts effectively act as downstream task experts, enabling specialization for new tasks without retraining the entire model.
Restricted Functional Expressiveness of Prompt Experts. Equation (4) reveals a key limitation of prompt experts $f_{N+1},\dots,f_{N+N_{p}}$. Although they are formally functions of $\bm{X}$}, they are represented by fixed prompt vectors ${\bm{p}}_{1},\dots,{\bm{p}}_{N_{p}}$ that remain constant regardless of the input. While their associated score functions $s_{i,N+j^{\prime}}$ in Equation (6) are input-dependent linear functions, the functional form realized by the prompt expert itself is static. This departs from typical MoE usage in the literature, where experts are adaptive functions of the input. Furthermore, it contrasts with the pre-trained experts $f_{1},\dots,f_{N}$, which are linear functions of $\bm{X}$} (see Equation (3)), making them comparatively more expressive. We hypothesize that this limited flexibility may constrain the effectiveness of VPT as a fine-tuning strategy. Supporting this, prior work Petrov et al. (2023) demonstrates that prompt tuning can only add a bias term to the output of an attention block, thereby restricting representational capacity. This observation naturally motivates the central question: Can the performance of visual prompt tuning be improved by making prompt experts more expressive? Addressing this, we propose a novel prompt formulation to enhance their expressiveness in Section 4.